A 9.450 mol sample of krypton gas is maintained in a 0.8100 L container at 300.1 K. What is the pressure in atm calculated using the van der Waals' equation for Kr gas under these conditions? For Kr, a = 2.318 L2atm/mol2 and b = 3.978×10-2 L/mol.

atm
Answer #1

Vander waal equation is given as,

\left ( P+\frac{an^2}{V^2} \right )(V-nb)=nRT\\ \Rightarrow P=\frac{nRT}{V-nb}-\frac{an^2}{V^2}

Here, P is the pressure, n = no. of moles = 9.450. V is Volume = 0.81 L, T is Temperature =300.1 K, a= 2.318 L2atm/mol2, b = 3.978 x 10-2 L/mol and R is the universal gas constant 0.082057 L atm mol-1 K-1

Putting in the values, we have,

P=\frac{(9.450 mol)(0.082057 L\cdot atm/mol\cdot K)(300.1 K)}{(0.8100L-(9.450 mol)(3.978\times10^{-2}L/mol))}-\frac{(2.318 L^2atm/mol^2)(9.450mol)^2}{(0.8100L)^2}\\ \Rightarrow P = 536.09 atm-315.5 atm=220.6 atm